OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

OpenStudy (amistre64):

what makes a sequence arithmetic?

OpenStudy (anonymous):

t14=a+13d ................i t15=a+14d.................ii solve these two equations first (simultaneous equations)

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

look, i'm solving it .observe it carefully,ok

OpenStudy (amistre64):

this is really simpler than solving a system of equation ... what makes a sequence arithmetic?

OpenStudy (anonymous):

I'm not really sure

OpenStudy (anonymous):

they have a common difference>?

OpenStudy (amistre64):

yes; a sequence is arithmetic of you can add the same value each time to get to the next number

OpenStudy (amistre64):

so, to get from the 14th to the 15th values, we added what? -33+d=9

OpenStudy (anonymous):

d=42?

OpenStudy (anonymous):

t14=a+13d ................i t15=a+14d.................ii so, now a+13d=-33......i a+14d=9..........ii now subtracting i from ii; -d=-42 d=42 now put d in equation ii; a+14*42=9 a=579 now tn=579+(n-1)42 =579+42n-42 =537+42n

OpenStudy (amistre64):

good then we have some setup such that: an = a1 + 42(n-1) using either of the given values we can easily determine a1

OpenStudy (amistre64):

a15 = a1 + 42(15-1) 9 = a1 + 42(14) 9 - 42(14)= a1

OpenStudy (anonymous):

a1 only has one value

OpenStudy (anonymous):

a1=-579?

OpenStudy (amistre64):

correct, so out equation is: an = -579 +42(n-1)

OpenStudy (anonymous):

yes lol i'm so sorry .yes a1=-579

OpenStudy (anonymous):

Awesome! thank you guys a lot!!

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

would u mind helping me with another?

OpenStudy (anonymous):

so answer will be -621+42n

OpenStudy (anonymous):

ur welcome.no not at all

OpenStudy (amistre64):

that result is suitable as well

OpenStudy (anonymous):

I'll post it and tag you guys

OpenStudy (anonymous):

thanx ami and jquelly for helping and correcting me

OpenStudy (anonymous):

ok